# SpiroLLM: Finetuning Pretrained LLMs to Understand Spirogram Time Series with Clinical Validation in COPD Reporting

### v1

@@ ABSTRACT
Chronic Obstructive Pulmonary Disease (COPD), a major chronic respiratory disease with persistent airflow limitation, is a leading global cause of disability and mortality. Respiratory spirogram time series, routinely collected during pulmonary function tests (PFTs), play a critical role in the early detection of repsiratory diseases and in monitoring lung function over time. However, most current AI models for COPD diagnosis are limited to outputting classification results without providing a rationale for their diagnostic process, while current Large Language Models (LLMs) cannot understand spirograms yet, which severely limits their clinical trust and adoption. To tackle this challenge, we leverage a cohort of 234,028 individuals from the UK Biobank (UKB) to propose SpiroLLM, the first multimodal large language model that can understand spirogram. The model extracts morphological features from respiratory curves via a SpiroEncoder and aligns them with PFT numerical values in a unified latent space using a SpiroProjector, ultimately empowering a large language model to generate a comprehensive diagnostic report. Experimental results confirm that SpiroLLM achieved a diagnostic AUROC of 0.8980 (95% CI: 0.8820-0.9132). In a robustness test with missing core data, it maintained a 100% valid response rate, far surpassing the 13.4% of a text-only model and showcasing the superiority of its multimodal design. This work demonstrates the substantial potential of deeply fusing physiological signals with large language models, establishing a new paradigm for the next generation of interpretable and reliable clinical decision support tools.
@@ INTRODUCTION
Chronic Obstructive Pulmonary Disease (COPD), a major chronic respiratory disease characterized by persistent airflow limitation, is one of the leading causes of disability and mortality worldwide1. In the clinical diagnosis and management of COPD, the Pulmonary Function Test (PFT) and its corresponding spirogram curve are indispensable. They not only represent the gold standard for diagnosis but also serve as a crucial basis for assessing disease severity, monitoring progression, and guiding treatment strategies2. However, the accurate interpretation of spirogram curves and the subsequent drafting of a standardized yet personalized diagnostic report are time-consuming, labor-intensive processes that are highly dependent on the specialized knowledge and long-term experience of clinicians. This reliance on expert resources is particularly pronounced in regions with limited medical access, creating a significant bottleneck in improving the efficiency and standardization of COPD diagnosis3.
To address this challenge, researchers have begun exploring the use of Artificial Intelligence (AI) to automate diagnostics4. In our prior work, we developed DeepSpiro 5, a model that demonstrated the feasibility of using deep learning to identify COPD-related features directly from spirogram curves. However, this and other early deep learning models were limited by their "black-box" nature, outputting only simple classification labels. Their inability to provide a rationale for their conclusions has hindered their clinical adoption and trust. More recently, the advent of Large Language Models (LLMs) has shown great promise in addressing this interpretability issue, with their ability to generate logically coherent medical texts that emulate the style of human experts 6. Nevertheless, applying LLMs to generate diagnostic reports directly from raw pulmonary function data still faces three core challenges:
@@ RESULTS
To comprehensively evaluate the performance of our proposed SpiroLLM model, we conducted a series of rigorous experiments and compared it against several key baseline models. These baselines include: 1) the base Llama3.1-8B large language model without any fine-tuning7; 2) the SpiroLLM-pftonly model, fine-tuned using only textualized pulmonary function metrics; and 3) a standalone DeepSpiro encoder model used solely for diagnostic classification 5. The evaluation primarily revolves around two core dimensions: diagnostic accuracy and the quality of the generated reports.
Diagnostic Accuracy
In terms of diagnostic accuracy, all fine-tuned models significantly outperformed the Llama3.1-8B baseline model and the standalone DeepSpiro classifier. As shown in Table 1, our multimodal model, SpiroLLM, achieved the best performance on both the AUROC (0.8980) and AUPRC (0.9049) metrics, demonstrating its superior classification capabilities.
@@
Table 1: Comparison of performance among different methods. Values are presented as mean (95% confidence interval).
Method Sensitivity Specificity F1 score AUROC AUPRC
DeepSpiro 5
0.6909
(0.6602-0.7219)
0.7898
(0.7646-0.8141)
0.7085
(0.6832-0.7331)
0.8266
(0.8077-0.8453)
0.8068
(0.7800-0.8326)
Llama3.1-8B 7
0.9842
(0.9753-0.9917)
0.1248
(0.1047-0.1458)
0.6424
(0.6207-0.6643)
0.7690
(0.7511-0.7863)
0.7318
(0.7097-0.7535)
SpiroLLM
0.7782
(0.7494-0.8058)
0.7574
(0.7305-0.7829)
0.7491
(0.7251-0.7709)
0.8965
(0.8810-0.9110)
0.9012
(0.8854-0.9154)
SpiroLLM
0.8327
(0.8071-0.8577)
0.6699
(0.6417-0.6992)
0.7435
(0.7208-0.7652)
0.8980
(0.8820-0.9132)
0.9049
(0.8890-0.9191)
Notably, the comparison between SpiroLLM and SpiroLLM-pftonly reveals that while the two are comparable in overall classification performance, SpiroLLM achieves a significantly higher diagnostic Sensitivity (0.8327 vs. 0.7782), representing a relative improvement of 7.0%. This result indicates that our proposed multimodal approach holds a distinct advantage in identifying true COPD patients, which is of great significance for reducing the rate of missed diagnoses in clinical practice.
@@
SpiroLLM (pftonly) 0.0048 (0.0000–0.0122) 0.5575 (0.5429–0.5726) 0.4790 (0.4549–0.5040)
0.7491 (0.7251–0.7709) 0.8965 (0.8810–0.9110) 0.9012 (0.8854–0.9154)
SpiroLLM 0.6990 (0.6761–0.7205) 0.8688 (0.8509–0.8862) 0.8193 (0.7931–0.8444)
0.7435 (0.7208–0.7652) 0.8980 (0.8820–0.9132) 0.9049 (0.8890–0.9191)
@@
The results of the expert evaluation are shown in Appendix E. Overall, the reports generated by SpiroLLM were of high quality. The experts generally agreed that a standout advantage of SpiroLLM is its precise interpretation and description of the pulmonary function curve morphology. They noted that SpiroLLM could accurately identify and describe typical visual features of COPD, such as a "concave descending limb," which provides strong evidence that our multimodal approach successfully translates visual signals into meaningful clinical descriptions. Furthermore, the experts affirmed the reliable clinical logic demonstrated by the model, noting its ability to correctly apply the GOLD standards to make key diagnostic judgments.
@@ DISCUSSION
This study has successfully designed, implemented, and validated a novel multimodal large language model framework named SpiroLLM, effectively addressing the challenge of automatically generating professional COPD diagnostic reports from raw PFT data. By combining a specially designed deep learning time-series encoder with the powerful reasoning and generative capabilities of large language models, this framework provides an end-to-end solution that is both accurate and robust. The experimental results strongly support the effectiveness of this approach, demonstrating that SpiroLLM significantly outperforms conventional single-modality methods in terms of diagnostic accuracy, quality of generated reports, and model robustness.
Our core contribution, SpiroLLM, is a novel multimodal framework that integrates a deep learning encoder with a large language model in an end-to-end fashion. This bridges the technical pathway from raw physiological time-series to the generation of professional diagnostic reports, addressing the fundamental problem of siloed model capabilities found in existing methods. To support the effective training of this framework, we further constructed a semi-automated gold-standard report generation pipeline. This pipeline, which combines multi-model collaboration and knowledge base augmentation, provides an effective and practical solution to the key bottleneck of scarce, high-quality annotated data that is prevalent in the medical AI field. Building on this, to scientifically measure the true value of our model, we designed a comprehensive evaluation scheme that goes beyond traditional text-overlap metrics. This scheme not only tailors evaluation criteria across six clinical dimensions for the "LLM-as-a-Judge" methodology but also innovatively introduces input masking experiments, allowing us to quantitatively confirm the independent contribution of multimodal fusion to enhancing model robustness.
The SpiroLLM framework proposed in this research holds significant potential for societal and clinical application. For clinicians, the model can serve as a powerful auxiliary diagnostic tool, significantly enhancing diagnostic efficiency by automatically generating high-quality draft reports. This frees clinicians from repetitive paperwork, allowing them to focus on judging complex cases and making critical decisions. Concurrently, the standardized report style can help improve diagnostic consistency across different tiers of medical institutions. From a public health perspective, an efficient and reliable automated diagnostic system can help expand the coverage of COPD screening, facilitating early identification and intervention, which in turn improves patient outcomes. Furthermore, our proposed semi-automated data annotation pipeline and comprehensive evaluation system also offer valuable methodological references for future medical AI research, holding broad applicability, especially in scenarios lacking large-scale manually annotated data.
Despite the encouraging results of this study, it is important to acknowledge several limitations. First, our model was trained and validated primarily on the UKB dataset, whose population is relatively homogeneous in terms of ethnicity. Therefore, the model’s generalization ability to other ethnic populations requires further validation. Second, while our evaluation is comprehensive, it relies primarily on retrospective data and automated evaluation metrics. Although the introduction of an LLM judge enhances the professionalism of the assessment, a gap may still exist between its results and the judgments of practicing physicians in a real clinical environment. Finally, the current model framework focuses on a single disease, COPD, and its applicability to other respiratory diseases has not yet been explored.
Future work will be centered around addressing these limitations and exploring broader applications. First, we will work to validate and optimize the model on datasets containing more diverse ethnic, geographic, and clinical characteristics to enhance its generalization ability. Second, we plan to deploy SpiroLLM in a simulated clinical environment and incorporate a feedback mechanism from real-world pulmonologists, forming a closed loop for continuous learning and iteration to further refine its diagnostic reasoning logic and report expression style. Finally, we will explore extending the framework to other respiratory diseases that also rely on pulmonary function tests, with the ultimate goal of developing it into a more general-purpose intelligent tool for pulmonary function interpretation and diagnostic report generation.

### v2

@@ ABSTRACT
Chronic Obstructive Pulmonary Disease (COPD), a major chronic respiratory disease with persistent airflow limitation, is a leading global cause of disability and mortality. Respiratory spirogram time series, routinely collected during pulmonary function tests (PFTs), play a critical role in the early detection of repsiratory diseases and in monitoring lung function over time. However, most current AI models for COPD diagnosis are limited to outputting classification results without providing a rationale for their diagnostic process, while current Large Language Models (LLMs) cannot understand spirograms yet, which severely limits their clinical trust and adoption. To tackle this challenge, we leverage a cohort of 234,028 individuals from the UK Biobank (UKB) to propose SpiroLLM, the first multimodal large language model that can understand spirogram. The model extracts morphological features from respiratory curves via a SpiroEncoder and aligns them with PFT numerical values in a unified latent space using a SpiroProjector, ultimately empowering a large language model to generate a comprehensive diagnostic report. Experimental results confirm that SpiroLLM achieved a diagnostic AUROC of 0.8977 (95% CI: 0.88-0.91). In a robustness test with missing core data, it maintained a 100% valid response rate, far surpassing the 13.4% of a text-only model and showcasing the superiority of its multimodal design. This work demonstrates the substantial potential of deeply fusing physiological signals with large language models, establishing a new paradigm for the next generation of interpretable and reliable clinical decision support tools.
@@ INTRODUCTION
Chronic Obstructive Pulmonary Disease (COPD), a major chronic respiratory disease characterized by persistent airflow limitation, is one of the leading causes of disability and mortality worldwideVenkatesan2025. In the clinical diagnosis and management of COPD, the Pulmonary Function Test (PFT) and its corresponding spirogram curve are indispensable. They not only represent the gold standard for diagnosis but also serve as a crucial basis for assessing disease severity, monitoring progression, and guiding treatment strategiesagusti2022global. However, the accurate interpretation of spirogram curves and the subsequent drafting of a standardized yet personalized diagnostic report are time-consuming, labor-intensive processes that are highly dependent on the specialized knowledge and long-term experience of clinicians. This reliance on expert resources is particularly pronounced in regions with limited medical access, creating a significant bottleneck in improving the efficiency and standardization of COPD diagnosisstanojevic2022ers.
To address this challenge, researchers have begun exploring the use of Artificial Intelligence (AI) to automate diagnosticsdas2023collaboration. In our prior work, we developed DeepSpiro mei2025deep, a model that demonstrated the feasibility of using deep learning to identify COPD-related features directly from spirogram curves. However, this and other early deep learning models were limited by their "black-box" nature, outputting only simple classification labels. Their inability to provide a rationale for their conclusions has hindered their clinical adoption and trust. More recently, the advent of Large Language Models (LLMs) has shown great promise in addressing this interpretability issue, with their ability to generate logically coherent medical texts that emulate the style of human experts liu2025application. Nevertheless, applying LLMs to generate diagnostic reports directly from raw pulmonary function data still faces three core challenges:
@@ RESULTS
Diagnostic Performance Analysis
We evaluated the diagnostic performance of SpiroLLM against the comprehensive set of baselines detailed in the Methods. Table 1 presents the comprehensive diagnostic performance comparison. Statistical significance was assessed using Bootstrap resampling ($n=1000$).
@@
Table 1: Comparison of performance among different methods. The table reports the Mean (95% Confidence Interval). The proposed SpiroLLM demonstrates superior performance, particularly in Sensitivity, compared to all baselines.
Category Model AUROC AUPRC Sensitivity Specificity F1 Score
Demographic & Clinical Baselines XGBoost-Demo chen2016xgboost 0.7288 (0.71-0.75) 0.6497 (0.61-0.69) 0.7082 (0.68-0.74) 0.6456 (0.62-0.67) 0.6597 (0.63-0.68)
XGBoost-Anthro chen2016xgboost 0.7368 (0.71-0.76) 0.6816 (0.65-0.71) 0.7703 (0.74-0.80) 0.5953 (0.57-0.62) 0.6783 (0.65-0.70)
XGBoost-Clinical chen2016xgboost 0.8573 (0.84-0.87) 0.8331 (0.81-0.86) 0.7377 (0.71-0.77) 0.8292 (0.81-0.85) 0.7569 (0.73-0.78)
Time-Series Baselines Informer zhou2021informer 0.7687 (0.75-0.79) 0.7098 (0.68-0.74) 0.7351 (0.70-0.77) 0.6800 (0.65-0.71) 0.6898 (0.66-0.71)
PatchTST Yuqietal-2023-PatchTST 0.8073 (0.79-0.83) 0.7810 (0.75-0.81) 0.7389 (0.71-0.77) 0.7378 (0.71-0.76) 0.7161 (0.69-0.74)
TimesNet wu2022timesnet 0.8270 (0.81-0.85) 0.8035 (0.78-0.83) 0.7582 (0.73-0.79) 0.7643 (0.74-0.79) 0.7397 (0.72-0.76)
SpiroEncoder mei2025deep 0.8249 (0.81-0.84) 0.8041 (0.78-0.83) 0.7474 (0.72-0.78) 0.7467 (0.72-0.77) 0.7253 (0.70-0.75)
ResNet18 cosentino2023inference 0.8355 (0.82-0.85) 0.8105 (0.78-0.84) 0.7401 (0.71-0.77) 0.7869 (0.76-0.81) 0.7386 (0.71-0.76)
Multimodal Fusion Baseline DeepSpiro mei2025deep 0.8625 (0.85-0.88) 0.8420 (0.82-0.87) 0.7838 (0.76-0.81) 0.7860 (0.76-0.81) 0.7651 (0.74-0.79)
Proposed Methods SpiroLLM (PFT-only) 0.8963†‡ (0.88-0.91) 0.9009†‡ (0.89-0.92) 0.7764 (0.75-0.80) 0.9960†‡ (0.99-1.00) 0.8717†‡ (0.85-0.89)
SpiroLLM 0.8977†‡ (0.88-0.91) 0.9046†‡ (0.89-0.92) 0.8165†‡∗ (0.79-0.84) 0.9931†‡ (0.99-1.00) 0.8947†‡∗ (0.88-0.91)
  • Notes:
  • Significant improvement over DeepSpiro ($p<0.001$).
  • Significant improvement over XGBoost-Clinical ($p<0.001$).
  • Significant improvement over SpiroLLM (PFT-only) ($p<0.001$).
  • Model Inputs Definition:
  • $\bullet$ XGBoost-Demo: Age, Sex, Smoking Status.
  • $\bullet$ XGBoost-Anthro: XGBoost-Demo features + Height, BMI.
  • $\bullet$ XGBoost-Clinical: XGBoost-Anthro features + PFT metrics derived from SpiroUtils (FEV1, FVC, etc.).
  • $\bullet$ Time-Series Baselines: Raw flow-volume curve data.
  • $\bullet$ DeepSpiro: Raw flow-volume curve data + Age, Sex, Smoking, and FEV1/FVC ratio.
  • $\bullet$ SpiroLLM (PFT-only): Basic Info (Age, Sex, Height, Smoking) + Comprehensive PFT Results (FEV1, FVC, FEV1/FVC, PEF, FEF25-75, including Measured, Predicted, LLN, and Z-scores).
  • $\bullet$ SpiroLLM: Fusion of SpiroLLM (PFT-only) inputs + Raw flow-volume curve data.
Consistent with clinical expectations, the progressive inclusion of comprehensive spirometry metrics effectively enhanced baseline performance. The XGBoost-Clinical model established a robust baseline, achieving an AUROC of 0.8573. SpiroLLM further elevated the AUROC to 0.8977, significantly outperforming this clinical baseline model ($p<0.001$). This result provides empirical evidence that raw flow-volume curves encapsulate subtle morphological patterns. These patterns, effectively captured by our visual encoder, offer complementary information to discrete pulmonary function test (PFT) metrics that cannot be fully characterized by numerical data alone.
Furthermore, SpiroLLM outperformed all unimodal time-series encoders (e.g., TimesNet, AUROC 0.8270) as well as the previously proposed multimodal benchmark, DeepSpiro (AUROC 0.8625, $p<0.001$). This underscores the architectural advantage of the proposed method: by aligning visual representations with the semantic reasoning capabilities of a Large Language Model (LLM), it achieves superior diagnostic discrimination compared to the traditional feature concatenation strategies employed in prior studies.
A critical comparison lies between the text-only variant (SpiroLLM PFT-only) and the complete SpiroLLM framework. Although their overall discriminative ability (AUROC) was comparable (0.8963 vs. 0.8977), the incorporation of raw flow-volume curves yielded a statistically significant improvement in sensitivity (81.65% vs. 77.64%, $p<0.001$). This absolute improvement of approximately 4.0% highlights the distinct clinical value of the visual modality: it enables the identification of "hard-to-diagnose" positive cases that present ambiguously in numerical metrics, thereby effectively reducing the rate of missed diagnoses in screening scenarios.
Visual Interpretability and Morphological Analysis
To investigate whether the observed performance enhancements—particularly the increased sensitivity within the multimodal framework—stem from clinically meaningful feature extraction rather than data artifacts, we visualized the attention distribution of the visual encoder (SpiroEncoder). By extracting activation maps from the final layer of the encoder and mapping them onto the raw flow-volume curves, we were able to interpret the specific regions prioritized by the model.
Figure 3: Interpretable visualization of the SpiroEncoder. Flow-volume curves are color-coded based on attention weights extracted from the final layer of the encoder (red indicates high attention; gray indicates low attention). (a-b) COPD cases: The model exhibits high attention toward the concave descending limb, successfully identifying the characteristic "scooped" pattern associated with airway obstruction. (c-d) Normal cases: Attention remains focused on the descending limb, validating the linear or convex profile associated with normal lung function. This confirms that the model explicitly learns to analyze morphological features indicative of small airway obstruction.
As illustrated in Figure 3, the generated heatmaps reveal a high degree of alignment between the features learned by the model and clinical expert intuition. Across all subjects, the model consistently concentrated regions of high attention (indicated in red) along the descending limb of the expiratory curve. In clinical physiology, this segment corresponds to variations in small airway airflow (FEF25–75%) and represents the primary site of airflow obstruction in Chronic Obstructive Pulmonary Disease (COPD). Specifically, in positive cases (Figures 3a and 3b), the model exhibited significant focus on the characteristic concavity or "scooped" pattern appearing after peak flow. Conversely, in normal cases (Figures 3c and 3d), attention was directed toward the relatively linear or convex trajectory of the descending limb.
These visualization results empirically demonstrate that the visual encoder serves as a critical component capable of autonomously learning to identify characteristic morphological landmarks of airflow limitation, thereby providing a transparent and interpretable basis for the model’s diagnostic decisions.
@@
SpiroLLM (PFT-only) 0.0685 (0.0000–0.1695) 0.5101 (0.4795–0.5468) 0.2273 (0.1721–0.2891)
0.8717 (0.8530–0.8892) 0.8963 (0.8811–0.9110) 0.9009 (0.8851–0.9154)
SpiroLLM 0.8486 (0.8298–0.8668) 0.8685 (0.8508–0.8856) 0.8186 (0.7926–0.8438)
0.8947 (0.8780–0.9107) 0.8977 (0.8814–0.9134) 0.9046 (0.8884–0.9197)
@@
The results of the expert evaluation are shown in Appendix F. Overall, the reports generated by SpiroLLM were of high quality. The experts generally agreed that a standout advantage of SpiroLLM is its precise interpretation and description of the pulmonary function curve morphology. They noted that SpiroLLM could accurately identify and describe typical visual features of COPD, such as a "concave descending limb," which provides strong evidence that our multimodal approach successfully translates visual signals into meaningful clinical descriptions. Furthermore, the experts affirmed the reliable clinical logic demonstrated by the model, noting its ability to correctly apply the GOLD standards to make key diagnostic judgments.
@@ DISCUSSION
In this study, we propose SpiroLLM, a multimodal framework that integrates raw flow-volume curves with clinical metadata. Evaluation using the UK Biobank cohort demonstrates a substantial improvement in diagnostic performance. First, the model constructed based on traditional clinical metrics (XGBoost-Clinical) established a robust baseline (AUROC 0.8573), corroborating the predictive efficacy of standard pulmonary function test (PFT) indicators. Second, by incorporating raw flow-volume curves, the existing multimodal baseline (DeepSpiro) achieved a marked improvement in sensitivity (78.38% vs. 73.77%), confirming that curve morphological information possesses independent diagnostic value. Most notably, SpiroLLM surpassed both baseline models, achieving a state-of-the-art AUROC of 0.8977 (p<0.001). This progression in performance indicates that while visual morphological data provides critical complementary information, our proposed method—which aligns visual features end-to-end with the semantic space of a Large Language Model (LLM)—more fully unlocks its diagnostic potential, significantly outperforming traditional feature concatenation strategies.
The performance enhancement stems primarily from the synergy between curve morphology analysis and the reasoning capabilities of the LLM. Although discrete PFT metrics possess high specificity, they often fail to capture subtle pathological patterns, resulting in a "sensitivity ceiling." The introduction of the visual encoder in DeepSpiro partially overcomes this limitation by identifying morphological anomalies missed by numerical thresholds. However, SpiroLLM achieves a far more significant breakthrough: compared to the text-only baseline, it improves sensitivity by an additional 4.0% (p<0.001). This suggests that SpiroLLM does not merely "recognize" curve morphology but effectively "interprets" its clinical significance, thereby bridging the gap between rigid numerical standards and expert intuition. This capability facilitates the identification of cases that remain elusive to simple fusion methods.
Beyond diagnostic accuracy, the reliability of this framework is grounded in two methodological innovations that ensure robustness in real-world scenarios. First, our adoption of a semi-automated pipeline for generating gold-standard reports offers a viable solution to the data scarcity issue prevalent in medical AI. By leveraging multi-model collaboration to synthesize high-quality training data, we demonstrate that effective supervisory signals can be obtained without incurring prohibitive manual annotation costs. Second, input masking experiments quantitatively validate the independent value of the visual modality: even when key quantitative metrics are removed, SpiroLLM maintains a high AUROC (0.8688) and a 100% effective response rate, whereas text-only variants exhibit systemic failure. This proves that our multimodal fusion is not merely redundant but constitutes a robust system capable of maintaining stability under conditions of incomplete data.
Translating these technical advantages into clinical practice holds significant implications for both individual patient care and public health. For clinicians, SpiroLLM serves as a potent decision-support tool, enhancing workflow efficiency by automatically generating high-quality, standardized reports. This allows physicians to be liberated from repetitive administrative tasks, enabling them to focus on clinical judgment for complex cases. From a broader public health perspective, such efficient and reliable automated systems promise to significantly expand screening coverage for Chronic Obstructive Pulmonary Disease (COPD) in resource-limited environments. By facilitating early identification and intervention, this framework provides a scalable pathway to improve long-term patient prognoses and mitigate the global burden of respiratory disease.
Despite these encouraging results, several limitations and ethical considerations warrant discussion. First, regarding data diversity, the model was trained primarily on the UK Biobank dataset, which is predominantly composed of white British participants; consequently, its generalizability to other ethnic groups remains to be verified. More importantly, the deployment of automated diagnostic systems carries inherent ethical risks, particularly regarding automation bias, which may be more acute in resource-limited settings lacking respiratory specialists. In such scenarios, overburdened clinicians might accept AI outputs without sufficient critical evaluation, effectively bypassing necessary clinical judgment. Therefore, SpiroLLM must be strictly defined as a "human-in-the-loop" decision-support tool rather than a fully autonomous diagnostic system. Furthermore, reliance on semi-automated data pipelines may introduce the risk of bias propagation—if the underlying knowledge base or the synthetic generation process reflects disparities in historical medical practice, the model may inadvertently amplify these biases. Consequently, continuous fairness monitoring and external validation are imperative prior to actual clinical application.
Future work will prioritize addressing the current limitations regarding demographic diversity and clinical validation. First, to enhance universality and algorithmic fairness, we plan to validate SpiroLLM across multi-center, multi-ethnic cohorts, expanding beyond the UK Biobank population to ensure robustness across diverse racial groups. Second, we intend to conduct prospective clinical pilot studies where SpiroLLM is integrated into real-world diagnostic workflows. This will allow us to quantitatively assess the model’s impact on clinical decision-making efficiency and report utility in a live clinical setting. In the long term, we aim to extend this framework to a broader spectrum of respiratory diseases, evolving SpiroLLM into a robust and versatile intelligent assistant for pulmonary medicine.